# Inexact Column Generation for Bayesian Network Structure Learning via Difference-of-Submodular Optimization

### v1

@@ 1 Introduction
Bayesian networks (BN) have wide applications in machine learning glymour1999computation . Its structure learning problem, i.e., identifying a directed acyclic graph (DAG) that represents the causal relationships between variables $(X_{i})_{i\in V}$ from data, is a central yet challenging topic in causal inference heinze2018causal . Existing methods for learning optimal Bayesian network structure from observational data fall into three main categories spirtes2000causation ; pearl2000models ; drton2017structure : (1) constraint-based approaches (e.g., PC spirtes2000causation ; colombo2014order and FCI spirtes2000causation ; zhang2008completeness ) that learn the structure through a constraint satisfaction problem and involve conditional independence tests to formulate the constraints; (2) score-based approaches (e.g., hill climbing heckerman1995learning , dynamic programming koivisto2004exact ; silander2012simple , integer programming cussens2020gobnilp and machine learning approaches yu2019dag ) that evaluate candidate DAGs using scoring functions and select the DAG structure with the highest score; (3) hybrid approaches that combine both score-based and constraint-based techniques (e.g., MMHC tsamardinos2006max ).
@@
Score-based BNSL is known to be NP-hard, even under the assumption that each node has a restricted number of parents chickering2004large . This inherent hardness of BNSL imposes a fundamental trade-off between solution quality and computational efficiency. Consequently, there are two lines of work that focus on the either side: (1) exact methods that guarantee optimality but often suffer from higher computational cost and (2) heuristic approaches that offer computational efficiency but with limited optimality guarantees. Within the BNSL literature, one of state-of-the-art exact approaches to identify the highest-scoring DAG is to formulate the problem as an Integer Program (IP) jaakkola2010learning ; 10.5555/3020548.3020567 ; chen2021integer , as represented by the GOBNILP method cussens2020gobnilp . While there are different ways to formulate the IP for this problem, a standard formulation involves binary variables that indicate whether to choose the specific set as the parent set for each node, and constraints that enforce the acyclicity requirement for the DAG. Given the observational data, the objective of the IP is to maximize the score, which is a function of the potential DAG structure measuring how well the data fits the model. Solving this IP is challenging since both the number of variables and the number of constraints grow exponentially with the node size of the DAG.
@@
Column Generation (CG) gilmore1961linear ; desaulniers2006column is a standard approach for addressing huge IPs as it effectively balances computational efficiency and solution quality. By dynamically generating variables as needed through pricing problems, CG can be used to derive efficient heuristic algorithms joncour2010column , or exact solution methods (often known as branch and price barnhart1998branch ) when combined with branch and cut. CG has been widely applied in operations research to solve real-world problems such as vehicle routing desrochers1992new and crew scheduling potthoff2010column . In cussens2012column ; cussens2023branch , Cussens introduces CG to BNSL to construct an exact branch and price algorithm. However, due to the complex pricing problems (which are formulated as mixed integer nonlinear programs in cussens2023branch ), the proposed approach has limited scalability and can only be applied to solve small BNSL instances.
@@
Indeed, to accelerate CG algorithms, it is advised that one should not try to solve all pricing problems exactly and an approximate solution often suffices until the last iteration (for proving optimality) lubbecke2005selected . In this paper, built on top of cussens2023branch , we advance CG solution techniques for BNSL. We develop an efficient pricing algorithm based on a reformulation of the pricing problem as a difference of submodular (DS) optimization problem, enabling efficient solution of the pricing problem via the Difference of Convex Algorithm (DCA) el2023difference ; iyer2012algorithms . Even though DCA does not necessarily provide global optimal solutions, it guarantees convergence and often generates high-quality candidate columns for our BNSL IP. Empirically, we demonstrate that it leads to effective primal heuristics, offering potential for developing more efficient exact solution approaches.
@@ 2.1 Scored-based BNSL as an IP
where $\texttt{score}(G;D)$ denotes the score of the graph $G=(V,E)$ under data $D$. Given a particular data type (continuous or discrete), there exist several different choices for the scoring function. In this paper, we focus on $\ell_{0}$-penalized likelihood scores, i.e., scoring functions of the form $\texttt{score}(G;D)=\log(L(G;D))-\Lambda\cdot k(G)$ for some $\Lambda\geq 0$, which favor graphs with a higher maximum likelihood $L(G;D)$ and a lower number of free parameters to be estimated $k(G)$ in the graphical model. By setting $\Lambda$ differently, one can recover some commonly used scoring functions, such as the AIC score akaike1974new and the BIC score schwarz1978estimating .
@@
The additive nature of the scoring function enables us to reformulate (1) as an integer (linear) program. A classic IP formulation based on cluster constraints 10.5555/3020548.3020567 ; jaakkola2010learning is as follows
@@
Here $\mathcal{P}_{i}:=2^{V\setminus\{i\}}$ denotes the set of candidate parent sets of node $i$, the binary decision variable $x_{i\leftarrow J}$ indicates whether parent set $J$ is selected for node $i$, with an associated local score $\texttt{score}_{i}(J)$. The linear objective reflects the score of the graph composed of the selected parent sets. Constraint (2b) enforces exactly one parent set to be chosen per node, while cluster constraints (2c) jaakkola2010learning guarantee acyclicity by preventing directed cycles among any nonempty node subset $C\in\mathcal{C}:=2^{V}\setminus\{\emptyset\}$.
Note that formulation (2) involves $\Theta(n2^{n-1})$ binary variables and $\Theta(2^{n})$ constraints. Due to its exponential size, even directly writing down (2) as an IP in a computer is infeasible for reasonably large $n$, not to say solving it. Although the exponential number of constraints can be addressed through row generation (i.e., replacing $\mathcal{C}$ by a small subset $\hat{\mathcal{C}}$ in (2) and iteratively adding violated ones into $\hat{\mathcal{C}}$, see Appendix B) 10.5555/3020548.3020567 , the exponential number of variables remains a significant challenge. To deal with exponentially many variables, existing BNSL IP approaches often rely on a (compromising) assumption that the size of the parent set for each node is no larger than some constant $k$, in which case one may replace $\mathcal{P}_{i}$ by $\{J\in\mathcal{P}_{i}:|J|\leq k\}$ in (2) for all $i$ to reduce the number of variables bartlett2017integer . Instead of putting such a restriction on the parent sets, we consider CG for BNSL as proposed in cussens2012column ; cussens2023branch .
@@ 2.2 Column Generation and the Pricing Problem for BNSL
Rather than including all the variables, CG handles IPs with huge number of variables by dynamically generating only the necessary ones through the so-called pricing problems. In the context of the BNSL IP (2), as demonstrated in cussens2023branch , the CG procedure begins with a restricted set of parent sets $\hat{\mathcal{P}}_{i}\subseteq\mathcal{P}_{i}$ for each node $i=1,\ldots,n$, and considers a restricted version of (2) only over variables $x_{i\leftarrow J},$ for $i=1,\ldots,n,~{}J\in\hat{\mathcal{P}}_{i}$, i.e., (2) with a small subset of columns. We refer to the linear programming (LP) relaxation of this restricted problem as the Restricted Master LP (RMLP). Based on the current RMLP solution, CG iteratively searches for additional parent set choices that may improve the objective value through the solution of a pricing problem for each node $i$, and add them to $\hat{\mathcal{P}}_{i}$. This pricing problem aims to optimize the so-called reduced costs, to identify high-quality parent set choices to gradually enlarge the variable space and solve the LP relaxation of (2).
As pointed out in cussens2023branch , the pricing problem for generating candidate parent sets for node $i$ can be formulated as the following optimization problem:
@@
where $\lambda_{i}^{*}$ and $\lambda_{C}^{*}\geq 0$ are the optimal dual values associated with constraints (2b) and (2c) in the RMLP, respectively. The value of $z_{i}(J;\lambda^{*})$ is known as the reduced cost bertsimas1997introduction of the variable $x_{i\leftarrow J}$, which measures the negative of how much the RMLP objective value might increase by adding the column associated with $x_{i\leftarrow J}$ to RMLP. The column generation process terminates when the optimal reduced cost (3) becomes nonnegative for all the nodes. In this paper, instead of using CG to construct an exact solution approach like branch and price, we consider a straightforward CG-based heuristic named the restricted master heuristic joncour2010column , where we solve the IP (2) with the candidate set $\hat{\mathcal{P}}_{i}$ generated by CG, namely the Restricted Master IP (RMIP), after CG (approximately) terminates.
@@ 2.3 Difference of Submodular Optimization via DCA
The pricing problem (3) is a set function optimization problem. Although problem (3), in the continuous Gaussian case, can be formulated as a mixed-integer nonlinear program (MINLP), solving it with an MINLP solver is highly computationally demanding as shown in cussens2023branch . Taking a direct set function optimization perspective, it is known that any set function can be written as a difference of submodular (DS) functions although finding such a DS decomposition has exponential complexity iyer2012algorithms . However, when such a decomposition is known, one can develop efficient algorithms to find local solutions iyer2012algorithms ; el2023difference . Here we briefly review a DS optimization algorithm el2023difference based on the well-known DCA an2005dc .
@@ A DS function $z:2^{V}\to\mathbb{R}$ is a set function that can be expressed as:
$z(J)=g(J)-f(J),$
where both $g:2^{V}\to\mathbb{R}$ and $f$ are submodular set functions (see Appendix A.2) fujishige2005submodular . For a set function $h:2^{V}\to\mathbb{R}$, its minimization can be equivalently formulated through its Lovász extension $h^{L}:[0,1]^{n}\to\mathbb{R}$ (see Appendix A.2) lovasz1983submodular . The convexity of the Lovász extension for submodular functions enables the reformulation of $\min_{J}z(J)=g(J)-f(J)$ as a difference of convex (DC) program
@@
As the subgradient of the Lovász extension of a submodular function can be efficiently evaluated through its function evaluation, the convex program (Line 5) in Algorithm 1 can be efficiently solved via bundle methods bagirov2014introduction . In our case, we adopt the classic Kelley’s Algorithm kelley1960cutting for numerical experiments.
@@
It is also easy to recover a solution of the original set function optimization problem from DCA, as the function value of $z$ agrees with $z^{L}$ at integer solutions and all extreme point solutions of the convex program (Line 5) in Algorithm 1 are integer solutions (since the corresponding set function is submodular).
@@ 3.1 The Pricing Problem as DS Optimization
A key step in CG is the solution of the pricing problem (3) that iteratively selects “promising” columns to add into the restricted problem formulation. Existing MINLP approaches are known to have very limited scalability to solve the pricing problem for BNSL cussens2023branch . We show in this section how the pricing problem can be explicitly rewritten as a DS optimization problem, enabling us to take advantage of the DCA algorithm we describe in Section 2.3.
@@
$g(J)=\log\operatorname{det}(\hat{\Sigma}_{J\cup i,J\cup i})+\Lambda|J|+\sum_{\begin{subarray}{c}C\in\hat{\mathcal{C}}:\\ i\in C,~{}J\cap C\neq\emptyset\end{subarray}}\lambda_{C}^{*}+\lambda_{i}^{*}$
@@
$g(J)=H(J\cup\{i\})+\sum_{\begin{subarray}{c}C\in\hat{\mathcal{C}}:\\ i\in C,~{}J\cap C\neq\emptyset\end{subarray}}\lambda_{C}^{*}+\lambda_{i}^{*}$
@@
$f(J)=\left(H(J)-\Lambda(a_{i}-1)\prod_{j\in J}a_{j}\right)$
@@ 3.2 Implementation of DCA for Pricing
To optimize this objective, we apply DCA, i.e., Algorithm  1, to iteratively minimize $z^{L}(x)$ through its successive convex approximations. The DCA procedure for solving the pricing problem for node $i$ proceeds as follows. The algorithm begins by initializing an $(n-1)$-dimensional solution vector $(x_{j})_{j\in V\setminus\{i\}}$. The required inputs include the dataset $D\in\mathbb{R}^{N\times n}$, node index $i$, optimal dual solutions $\lambda^{*}$ of RMLP, regularization parameter $\Lambda$, and convergence thresholds $\epsilon$.
@@
$y^{t}_{\sigma(k)}=\begin{cases}\log\text{det}(\hat{\Sigma}_{\{\sigma(1)\}})&k=1\\ \log\text{det}(\hat{\Sigma}_{\{\sigma(1),\ldots,\sigma(k)\}})-\log\text{det}(\hat{\Sigma}_{\{\sigma(1),\ldots,\sigma(k-1)\}})&k\geq 2,\end{cases}$
where $\hat{\Sigma}_{\{\sigma(1),\ldots,\sigma(k)\}}$ denotes the empirical covariance matrix for the corresponding columns $\{\sigma(1),\ldots,\sigma(k)\}$ of the data matrix $D\in\mathbb{R}^{N\times n}$. For the discrete case, a subgradient $y^{t}$ of $f^{L}$ at $x^{t}$ is given by
@@
After evaluating $\log\text{det}(\hat{\Sigma}_{({\cdot})})$ or $H(\cdot)$ during the computation of $y^{t}$, it is crucial to save the evaluated values to avoid repeating calculations in the future. This small trick greatly improves the efficiency of DCA in pricing problem solutions in practice.

### v2

@@ 1 Introduction
Bayesian networks (BN) have wide applications in machine learning [27]. Its structure learning problem, i.e., identifying a directed acyclic graph (DAG) that represents the causal relationships between variables $(X_{i})_{i\in V}$ from data, is a central yet challenging topic in causal inference [29]. Existing methods for learning optimal Bayesian network structure from observational data fall into three main categories [45, 40, 21]: (1) constraint-based approaches (e.g., PC [45, 14] and FCI [45, 50]) that learn the structure through a constraint satisfaction problem and involve conditional independence tests to formulate the constraints; (2) score-based approaches (e.g., hill climbing [28], dynamic programming [35, 44], integer programming [17] and machine learning approaches [49]) that evaluate candidate DAGs using scoring functions and select the DAG structure with the highest score; (3) hybrid approaches that combine both score-based and constraint-based techniques (e.g., MMHC [47]).
@@
Score-based BNSL is known to be NP-hard, even under the assumption that each node has a restricted number of parents [13]. This inherent hardness of BNSL imposes a fundamental trade-off between solution quality and computational efficiency. Consequently, there are two lines of work that focus on the either side: (1) exact methods that guarantee optimality but often suffer from higher computational cost and (2) heuristic approaches that offer computational efficiency but with limited optimality guarantees. Within the BNSL literature, one of state-of-the-art exact approaches to identify the highest-scoring DAG is to formulate the problem as an Integer Program (IP) [31, 15, 12], as represented by the GOBNILP method [17]. While there are different ways to formulate the IP for this problem, a standard formulation involves binary variables that indicate whether to choose the specific set as the parent set for each node, and constraints that enforce the acyclicity requirement for the DAG. Given the observational data, the objective of the IP is to maximize the score, which is a function of the potential DAG structure measuring how well the data fits the model. Solving this IP is challenging since both the number of variables and the number of constraints grow exponentially with the node size of the DAG.
@@
Column Generation (CG) [26, 19] is a standard approach for addressing huge IPs as it effectively balances computational efficiency and solution quality. By dynamically generating variables as needed through pricing problems, CG can be used to derive efficient heuristic algorithms [32], or exact solution methods (often known as branch and price [5]) when combined with branch and cut. CG has been widely applied in operations research to solve real-world problems such as vehicle routing [20] and crew scheduling [41]. In [16, 18], Cussens introduces CG to BNSL to construct an exact branch and price algorithm. However, due to the complex pricing problems (which are formulated as mixed integer nonlinear programs in [18]), the proposed approach has limited scalability and can only be applied to solve small BNSL instances.
@@
Indeed, to accelerate CG algorithms, it is advised that one should not try to solve all pricing problems exactly and an approximate solution often suffices until the last few iterations (for proving optimality) [37]. In this paper, built on top of [18], we advance CG solution techniques for BNSL. We develop an efficient pricing algorithm based on a reformulation of the pricing problem as a difference of submodular (DS) optimization problem, enabling efficient solution of the pricing problem via the Difference of Convex Algorithm (DCA) [22, 30]. Even though DCA does not necessarily provide global optimal solutions, it guarantees convergence and often generates high-quality candidate columns for our BNSL IP. Empirically, we demonstrate that it leads to effective primal heuristics, offering potential for developing more efficient exact solution approaches.
@@ 2.1 Scored-based BNSL as an IP
where $\texttt{score}(G;D)$ denotes the score of the graph $G=(V,E)$ under data $D$ (see Appendix A.1). Given a particular data type (continuous or discrete), there exist several different choices for the scoring function. In this paper, we focus on $\ell_{0}$-penalized likelihood scores, i.e., scoring functions of the form $\texttt{score}(G;D)=\log(L(G;D))-\Lambda\cdot k(G)$ for some $\Lambda\geq 0$, which favor graphs with a higher maximum likelihood $L(G;D)$ and a lower number of free parameters to be estimated $k(G)$ in the graphical model. By setting $\Lambda$ differently, one can recover some commonly used scoring functions, such as the AIC score [1] and the BIC score [42].
@@
The additive nature of the scoring function enables us to reformulate (1) as an integer (linear) program. A classic IP formulation based on cluster constraints [15, 31] is as follows
@@
Here $\mathcal{P}_{i}:=2^{V\setminus\{i\}}$ denotes the set of candidate parent sets of node $i$, the binary decision variable $x_{i\leftarrow J}$ indicates whether parent set $J$ is selected for node $i$, with an associated local score $\texttt{score}_{i}(J)$. The linear objective reflects the score of the graph composed of the selected parent sets. Constraint (2b) enforces exactly one parent set to be chosen per node, while cluster constraints (2c) [31] guarantee acyclicity by preventing directed cycles among any nonempty node subset $C\in\mathcal{C}:=2^{V}\setminus\{\emptyset\}$.
Note that formulation (2) involves $\Theta(n2^{n-1})$ binary variables and $\Theta(2^{n})$ constraints. Due to its exponential size, even directly writing down (2) as an IP in a computer is infeasible for reasonably large $n$, not to say solving it. Although the exponential number of constraints can be addressed through row generation (i.e., replacing $\mathcal{C}$ by a small subset $\hat{\mathcal{C}}$ in (2) and iteratively adding violated ones into $\hat{\mathcal{C}}$, see Appendix B) [15], the exponential number of variables remains a significant challenge. To deal with exponentially many variables, existing BNSL IP approaches often rely on a (compromising) assumption that the size of the parent set for each node is no larger than some constant $k$, in which case one may replace $\mathcal{P}_{i}$ by $\{J\in\mathcal{P}_{i}:|J|\leq k\}$ in (2) for all $i$ to reduce the number of variables [6]. Instead of putting such a restriction on the parent sets, we consider CG for BNSL as proposed in [16, 18].
@@ 2.2 Column Generation and the Pricing Problem for BNSL
Rather than including all the variables, CG handles IPs with huge number of variables by dynamically generating only the necessary ones through the so-called pricing problems. In the context of the BNSL IP (2), as demonstrated in [18], the CG procedure begins with a restricted set of parent sets $\hat{\mathcal{P}}_{i}\subseteq\mathcal{P}_{i}$ for each node $i=1,\ldots,n$, and considers a restricted version of (2) only over variables $x_{i\leftarrow J},$ for $i=1,\ldots,n,~J\in\hat{\mathcal{P}}_{i}$, i.e., (2) with a small subset of columns. We refer to the linear programming (LP) relaxation of this restricted problem as the Restricted Master LP (RMLP). Based on the current RMLP solution, CG iteratively searches for additional parent set choices that may improve the objective value through the solution of a pricing problem for each node $i$, and add them to $\hat{\mathcal{P}}_{i}$. This pricing problem aims to optimize the so-called reduced costs, to identify high-quality parent set choices to gradually enlarge the variable space and solve the LP relaxation of (2).
As pointed out in [18], the pricing problem for generating candidate parent sets for node $i$ can be formulated as the following optimization problem:
@@
where $\lambda_{i}^{*}$ and $\lambda_{C}^{*}\geq 0$ are the optimal dual values associated with constraints (2b) and (2c) in the RMLP, respectively. The value of $z_{i}(J;\lambda^{*})$ is known as the reduced cost [8] of the variable $x_{i\leftarrow J}$. Its negative measures how much the RMLP objective value might increase by adding the column associated with $x_{i\leftarrow J}$ to RMLP. The column generation process terminates when the optimal reduced cost (3) becomes nonnegative for all the nodes. In this paper, instead of using CG to construct an exact solution approach like branch and price, we consider a straightforward CG-based heuristic named the restricted master heuristic [32], where we solve the IP (2) with the candidate set $\hat{\mathcal{P}}_{i}$ generated by CG, namely the Restricted Master IP (RMIP), after CG (approximately) terminates.
@@ 2.3 Difference of Submodular Optimization via DCA
The pricing problem (3) is a set function optimization problem. Although problem (3), in the continuous Gaussian case, can be formulated as a mixed-integer nonlinear program (MINLP), solving it with an MINLP solver is highly computationally demanding as shown in [18]. Taking a direct set function optimization perspective, it is known that any set function can be written as a difference of submodular (DS) functions although finding such a DS decomposition has exponential complexity [30]. However, when such a decomposition is known, one can develop efficient algorithms to find local solutions [30, 22]. Here we briefly review a DS optimization algorithm [22] based on the well-known DCA [2].
@@ A DS function $z:2^{V}\to\mathbb{R}$ is a set function that can be expressed as:
z(J)=g(J)f(J),z(J)=g(J)-f(J),
where both $g:2^{V}\to\mathbb{R}$ and $f$ are submodular set functions (see Appendix A.2) [25]. For a set function $h:2^{V}\to\mathbb{R}$, its minimization can be equivalently formulated through its Lovász extension $h^{L}:[0,1]^{n}\to\mathbb{R}$ (see Appendix A.2) [36]. The Lovász extensions for submodular functions are convex [36], enabling the reformulation of $\min_{J}z(J)=g(J)-f(J)$ as a difference of convex (DC) program
@@
As the subgradient of the Lovász extension of a submodular function can be efficiently evaluated through its function evaluation, the convex program (Line 5) in Algorithm 1 can be efficiently solved via bundle methods [4]. In our case, we adopt the classic Kelley’s Algorithm [33] for numerical experiments.
@@
It is also easy to recover a solution of the original set function optimization problem from DCA, as the function value of $z$ agrees with $z^{L}$ at integer solutions. Since the convex function $g^{L}(x)-\langle y^{t},x\rangle$ is the Lovász extension (which characterizes the convex envelope) of a submodular function (i.e., $g$ plus a modular function), all extreme point solutions of the convex program (Line 5) in Algorithm 1 are integer solutions.
@@ 3.1 The Pricing Problem as DS Optimization
A key step in CG is the solution of the pricing problem (3) that iteratively selects “promising” columns to add into the restricted problem formulation. Existing MINLP approaches are known to have very limited scalability to solve the pricing problem for BNSL [18]. We show in this section how the pricing problem can be explicitly rewritten as a DS optimization problem, enabling us to take advantage of the DCA algorithm we describe in Section 2.3.
@@
$g(J)=\frac{N}{2}\log\operatorname{det}(\hat{\Sigma}_{J\cup i,J\cup i})+\Lambda|J|+\sum_{\begin{subarray}{c}C\in\hat{\mathcal{C}}:\\ i\in C,~J\cap C\neq\emptyset\end{subarray}}\lambda_{C}^{*}+\lambda_{i}^{*}$
@@
$g(J)=N\cdot H(J\cup\{i\})+\sum_{\begin{subarray}{c}C\in\hat{\mathcal{C}}:\\ i\in C,~J\cap C\neq\emptyset\end{subarray}}\lambda_{C}^{*}+\lambda_{i}^{*}$
@@
$f(J)=N\cdot H(J)-\Lambda(a_{i}-1)\prod_{j\in J}a_{j}$
@@ 3.2 Implementation of DCA for Pricing
To optimize this objective, we apply DCA, i.e., Algorithm  1, to iteratively minimize $z^{L}(x)$ through its successive convex approximations. The DCA procedure for solving the pricing problem for node $i$ proceeds as follows. The algorithm begins by initializing an (n1)(n-1)-dimensional solution vector $(x_{j})_{j\in V\setminus\{i\}}$. The required inputs include the dataset $D\in\mathbb{R}^{N\times n}$, node index $i$, optimal dual solutions $\lambda^{*}$ of RMLP, regularization parameter $\Lambda$, and convergence thresholds $\epsilon$.
@@
[…]
Lemma 1.
Let $M$ be a $p\times p$ positive definite symmetric matrix with lower-triangular Cholesky factor $L$. In block form, these matrices can be written as
$M=\begin{bmatrix}M_{11}&M_{12}\\ M_{21}&M_{22}\end{bmatrix},\quad L=\begin{bmatrix}L_{1}&O\\ Y&L_{2}\end{bmatrix}$,
where $M_{11}$ and $L_{1}$ are $m\times m$ blocks ($m<p$) and $O$ is a zero matrix of appropriate size. Then $L_{1}$ is the Cholesky factor of $M_{11}$, and consequently,
$\log\operatorname{det}(M_{11})=\log(\operatorname{det}(L_{1})^{2})=2\sum_{k=1}^{m}\log L_{k,k}.$
For the discrete case, a subgradient $y^{t}$ of $f^{L}$ at $x^{t}$ is given by
@@
[…]